OpenStudy (anonymous):

What is the solution to the equation x + 9.5 = 27.5?

OpenStudy (usukidoll):

subtract 9.5 on both sides. remember to line up the decimal points.

OpenStudy (decentnabeel):

\[x+9.5=27.5\quad :\quad x=18\]

pooja195 (pooja195):

@DecentNabeel better to guide than give the answer

OpenStudy (decentnabeel):

ok @pooja195

OpenStudy (anonymous):

thx i accidentally put this question in

OpenStudy (usukidoll):

it's alright . nice username though @3.141592653589

OpenStudy (anonymous):

thx

OpenStudy (decentnabeel):

\[\mathrm{Multiply\:both\:sides\:by\:}10\] \[10x+10\cdot \:9.5=10\cdot \:27.5\] \[10x+95=275\] \[\mathrm{Subtract\:}95\mathrm{\:from\:both\:sides}\] \[10x+95-95=275-95\] \[10x=180\] divided 10 on both sides x=18